A solution is prepared at 25C that is initially 0.50 M in chlorous acid (HClO2), a weak acid with Ka= 1.1 x 10^-2, and 0.4M in (NaClO2). Calc the pH of the solution. Round 2 decimal places
Answer #1

Ka = 1.1*10^-2

pKa = - log (Ka)

= - log(1.1*10^-2)

= 1.959

use:

pH = pKa + log {[conjugate base]/[acid]}

= 1.959+ log {0.4/0.5}

= 1.862

Answer: 1.86